# Hybrid Attention Network: An efficient approach for anatomy-free landmark detection

### v1

@@ Abstract
Accurate anatomical landmark detection in medical images is crucial for clinical applications. Existing methods often struggle to balance global context with computational efficiency, particularly with high-resolution images. This paper introduces the Hybrid Attention Network(HAN), a novel hybrid architecture integrating CNNs and Transformers. Its core is the BiFormer module, utilizing Bi-Level Routing Attention (BRA) for efficient attention to relevant image regions. This, combined with Convolutional Attention Blocks (CAB) enhanced by CBAM, enables precise local feature refinement guided by the global context. A Feature Fusion Correction Module (FFCM) integrates multi-scale features, mitigating resolution loss. Deep supervision with MSE loss on multi-resolution heatmaps optimizes the model. Experiments on five diverse datasets demonstrate state-of-the-art performance, surpassing existing methods in accuracy, robustness, and efficiency. The HAN provides a promising solution for accurate and efficient anatomical landmark detection in complex medical images. Our codes and data will be released soon at: https://github.com/MIRACLE-Center/.
@@ 1 Introduction
Anatomical landmark detection is a core task in medical image analysis, widely used in clinical diagnosis, surgical planning, and treatment evaluation [1, 2]. Accurate landmark localization provides reliable anatomical references for clinicians, enabling automated diagnosis and personalized treatment [3]. It also supports foundational tasks such as image registration [4], segmentation [5], and 3D reconstruction [6], improving the accuracy of these tasks and advancing medical imaging technology. However, complex anatomical structures, individual variability, low signal-to-noise ratio and resolution differences in medical images make automated landmark detection challenging.
@@
Early landmark detection methods rely on handcrafted image features [7, 8, 9]. While these methods achieve some success in specific applications, they depend heavily on data quality and feature design. As a result, they struggle to handle complex anatomical structures and adapt to varying imaging conditions, leading to limited robustness. With the rapid advancement of deep learning, Convolutional Neural Networks (CNNs) have transformed image processing tasks, particularly in anatomical landmark detection and medical image segmentation. Models such as U-Net [10] and its variants [11, 12, 13, 14, 15] are now widely applied for landmark detection. These models employ encoder-decoder architectures to effectively integrate low-level features with high-level semantic information.
@@
Transformer architectures, with their self-attention mechanisms, are particularly effective at modeling global dependencies and capturing long-range spatial information. In the field of medical imaging, Transformers have been increasingly integrated into image analysis tasks, often in combination with convolution-based U-Net models such as UTNet [16], TransUNet [17], and nnFormer [18]. Since accurate landmark detection relies heavily on global context, incorporating Transformers to capture this information has been shown to enhance detection accuracy. Notable examples include models like SpineHRformer [19] and DATR [20].
@@
To leverage the complementary strengths of Convolutional Neural Networks (CNNs) and Transformers, hybrid architectures such as HTC [21] and CephalFormer [22] have emerged. These models combine CNNs for local feature extraction and Transformers for modeling global dependencies, achieving significant success in anatomical landmark detection. However, the computational complexity of standard self-attention mechanisms scales quadratically with the number of tokens, which limits the computational efficiency and overall performance of the model to some extent. Some methods, such as Gated Axial UNet (MedT) [23], have introduced sparsity to alleviate the computational burden, but they rely on manually designed, fixed sparse patterns. These fixed patterns limit the model’s adaptability when processing different queries, as they cannot dynamically adjust the sparse pattern according to the specific context. Medical images are typically characterized by high complexity and low signal-to-noise ratios, and there is a significant domain gap between datasets from different anatomical regions and acquisition protocols. These characteristics demand that models be able to flexibly select and adjust computation regions to address these challenges. Therefore, developing a sparse attention mechanism that dynamically adjusts based on queries tailored for medical image landmark detection tasks remains one of the key research directions to improve detection accuracy.
@@
To leverage the complementary strengths of Convolutional Neural To address the challenges in medical image landmark detection and improve accuracy and robustness, we propose a novel hybrid model incorporating dynamic sparse attention into a Transformer-CNN architecture. The method combines Convolutional Attention Blocks (CAB) with the BiFormer module, leveraging bilevel routing attention to effectively model long-range spatial dependencies while optimizing computational and memory efficiency. Specifically, the CAB module enhances feature selection through channel and spatial attention mechanisms, while the BiFormer module captures long-range dependencies via multi-scale feature extraction and self-attention, improving feature representation. Additionally, a Feature Fusion Correction Module(FFCM) is introduced to integrate global and local features, further enhancing the accuracy of landmark detection in medical images.
@@
We conduct extensive experiments across five public datasets, encompassing various anatomical regions, imaging methods, and resolutions, consistently achieving state-of-the-art (SOTA) results. For example, on the ISBI2015 dataset, our method achieves an MRE of 1.13mm, improving by 5% compared to HTC [21] (1.19mm) and 11% compared to FARNet [15] (1.27mm). Additionally, the Success Detection Rate (SDR) at 2mm reaches 84.78%, surpassing HTC by 1.54% and FARNet by 2.27%. Similar improvements are observed across all datasets, demonstrating the robustness, adaptability, and clear superiority of our method in high-resolution medical image analysis and landmark detection tasks.
@@
  • We propose a hybrid Transformer-CNN architecture that combines BiFormer modules and CAB to enhance feature selectivity and capture global dependencies, improving the accuracy and robustness of anatomical landmark detection.
  • We design a Feature Fusion Correction Module (FFCM) that effectively integrates global and local information, enabling the model to adapt to diverse anatomical regions and imaging conditions.
@@ 2 Related Work
2.1 Traditional Approaches to Landmark Detection
@@
Although these methods achieve some success in the early stages of landmark detection, they gradually fall short due to the high complexity and precision requirements of medical images. With the rapid development of deep learning, researchers begin exploring new approaches to overcome the limitations of traditional methods, significantly improving the performance and adaptability of landmark detection.
2.2 CNNs and Transformer Based Methods
With the rapid advancements in deep learning, Convolutional Neural Networks (CNNs) have been widely adopted for anatomical landmark detection. For instance, O’Neil et al. [13] proposes a two-stage method based on Fully Convolutional Networks (FCN). Payer et al. [14] introduces Spatial Configuration Network (SCN), which combines local appearance features with spatial configuration information. Zhu et al. [33] proposes GU2Net, a general model for multi-domain landmark detection, which improves robustness through cross-domain learning. Additionally, Ao et al. [15] develops FARNet, which incorporates multi-scale feature aggregation and refinement modules, further improving detection precision.
@@ 2.3 CNN-Transformer Hybrid Models
However, due to the complexity and low signal-to-noise ratio of medical images, high-performance models are necessary. The computational complexity of standard Transformer models becomes a bottleneck, especially when processing high-resolution medical images. Some studies, such as Swin-Unet [35] and Gated Axial UNet (MedT) [23], have incorporated sparsity into their models to address this challenge. Swin-CE [36] integrates Swin Transformer encoders with convolutional encoders for cephalometric landmark detection, but these methods often rely on manually designed sparse patterns, which lack flexibility in adapting to dynamic queries.
Recent advancements in vision Transformers, particularly those utilizing dynamic sparsity, provide promising solutions. For instance, dynamic token sparsity mechanisms prune a large number of non-informative tokens, thereby accelerating the model while maintaining accuracy. In [37], dynamic token sparsity was applied to reduce computational load by removing non-informative tokens without sacrificing precision. The application of dynamic and query-aware sparse attention mechanisms for landmark detection in medical images presents a significant research opportunity. In this study, we propose employing bi-level routing attention as a foundational dynamic sparse block, constructing a U-shaped hybrid CNN-Transformer encoder-decoder architecture for efficient landmark detection in medical images.
@@ 3 Proposed Methods
Figure 1: The overview of proposed Hybrid Attention Network(HAN). BiFormer is a module based on bilevel routing attention, and CAB stands for Convolutional Attention Blocks. Further details will be discussed later.
@@ 3.1 Problem Definition
The task of anatomical landmark detection aims to accurately predict the locations of multiple anatomical landmarks from input medical images. Formally, given an input image $I\in\mathbb{R}^{H\times W\times C}$, the objective is to predict a set of landmarks $L=\{l_{1},l_{2},\dots,l_{N}\}$, where each $l_{i}=(x_{i},y_{i})$ represents the position of a landmark in the image. $N$ represents the number of landmarks in the image, which is also the number of output heatmap channels. The corresponding heatmap $H_{i}(x_{i},y_{i})$ for each landmark $l_{i}$ is generated using a Gaussian function:
$H_{i}=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma^{2}}\right)$ (1)
where $C$ represents the number of image channels, with $C=1$ for X-ray images. $H$ and $W$ represent the image height and width, respectively. $\sigma$ represents the standard deviation of the Gaussian function used to generate the heatmap.
@@ 3.2 BiFormer Module Design
The BiFormer module, based on the Bi-Level Routing Attention (BRA) mechanism [37], is a key component of the proposed architecture. BRA is a dynamic, query-aware sparse attention mechanism designed to prune the least relevant key-value patchs at a coarse level, retaining only the most relevant parts for fine-grained token-level attention. Token-to-token attention is then computed within these selected patchs. Compared to static sparse attention mechanisms [35, 38], BRA is more flexible in modeling long-range dependencies, making it particularly effective for dense prediction tasks that require capturing global features.
@@
Furthermore, as shown in Fig. 3, to compute the token-to-token attention for a patch $i$, initially, collect the top-$k$ most relevant patch index from matrix $I^{p}$ of patch $i$; subsequently, apply the gather algorithm to extract the corresponding $K^{g}$ and $V^{g}$ values of the relevant patchs, where $K^{g},V^{g}\in\mathbb{R}^{S^{2}\times k\frac{HW}{S^{2}}\times C}$. Finally, calculate the attention between patch $i$ and the top-$k$ most relevant patchs, incorporating the Local Context Enhancement (LCE), to obtain the final output $O$ for this patch. The above process can be formulated as follows:
@@
Figure 2: (a) Architecture of the Proposed BiFormer Block. (b) Architecture of the Proposed Convolutional Attention Block. (c) Overview of the Convolutional Block Attention Module (CBAM).
@@ 3.3 Overal Network Architecture
As demonstrated in Fig. 1, the overall HAN employs a U-shaped architecture, with the main backbone consisting of hybrid Transformer-CNN modules, upsampling layers, and residual connections. The hybrid Transformer-CNN module is composed of BiFormer and CAB connected in series. The global features extracted by BiFormer are fed into CAB, which further refines local features, achieving an effective integration of global and local features. Specifically, the global information provided by BiFormer helps the model understand the geometric relationships between landmarks, while CAB refines local features based on these global features, making landmark detection more precise.
Compared to the original CNN, CAB is integrated with an attention module Convolutional Block Attention Module(CBAM) to select representative features. As demonstrated in Fig. 2(c), the CBAM module combines channel attention and spatial attention, allowing the network to focus more effectively on task-relevant features, enhancing the effectiveness of feature representation and the overall performance of the network. The role of CBAM can be formulated as:
$F_{\text{out}}=\text{SAM}(\text{CAM}(F_{\text{in}})\odot F_{\text{in}})\odot(\text{CAM}(F_{\text{in}})\odot F_{\text{in}}),$
where CAM represents channel attention, SAM represents spatial attention, and $\odot$ denotes element-wise multiplication.
As for input features, CAB first processes them using a dilated convolution with dilation set to 1, followed by normalization and activation function processing. Then, the processed features are passed through another dilated convolution with dilation set to 2. After normalization, the features are passed through the CBAM, and finally, the output is obtained through a residual connection. The above process illustrated in Fig. 2(b) and can be represented as:
@@
BiFormer provides rich contextual information to CAB, reducing redundant features and information loss, thereby accelerating the training process and improving convergence speed. Therefore, combining BiFormer with CAB not only enhances the model’s performance in landmark detection but also improves training efficiency and model robustness, making it suitable for complex medical images and other scenarios that require precise modeling.
As Fig. 1 shows, the downsampling part of the model consists of five layers of hybrid Transformer-CNN modules. For an input image, Patch embedding is first performed, followed by the BiFormer module for capturing global features and the CAB module for extracting local features. Multi-scale features $\{D_{1},D_{2},D_{3},D_{4},D_{5}\}$ are generated to enhance the model’s feature representation capability. These features are subsequently mapped to a feature space with 256 channels.
@@ 4.2 Baseline Methods
To verify the effectiveness of our proposed model, we compare it with several representative or SOTA methods, including U-Net [10], SCN [14], GU2Net [33], FARNet [15], DATR [20], SR-UNet , HTC [21], and CeLDA [41]:
@@
  • SRLD-Net[44]: Applies super-resolution networks and pyramid pooling to reduce errors and improve detection precision.

### v2

# HYATT-Net is Grand: A Hybrid Attention Network for Performant Anatomical Landmark Detection

@@ Abstract
Anatomical landmark detection (ALD) from a medical image is crucial for a wide array of clinical applications. While existing methods achieve quite some success in ALD, they often struggle to balance global context with computational efficiency, particularly with high-resolution images, thereby leading to the rise of a natural question: where is the performance limit of ALD? In this paper, we aim to forge performant ALD by proposing a HYbrid ATTention Network (HYATT-Net) with the following designs: (i) A novel hybrid architecture that integrates CNNs and Transformers. Its core is the BiFormer module, utilizing Bi-Level Routing Attention for efficient attention to relevant image regions. This, combined with Attention Residual Module(ARM), enables precise local feature refinement guided by the global context. (ii) A Feature Fusion Correction Module that aggregates multi-scale features and thus mitigates a resolution loss. Deep supervision with a mean-square error loss on multi-resolution heatmaps optimizes the model. Experiments on five diverse datasets demonstrate state-of-the-art performance, surpassing existing methods in accuracy, robustness, and efficiency. The HYATT-Net provides a promising solution for accurate and efficient ALD in complex medical images. Our codes and data are already released at: https://github.com/ECNUACRush/HYATT-Net.
@@ 1 Introduction
Anatomical landmark detection (ALD) is a core task in medical image analysis, widely used in clinical diagnosis, surgical planning, and treatment evaluation [1, 2]. Accurate landmark localization provides reliable anatomical references for clinicians, enabling automated diagnosis and personalized treatment [3]. It also supports foundational tasks such as image registration [4], segmentation [5, 6], and 3D reconstruction [7], improving the accuracy of these tasks and advancing medical imaging technology. However, complex anatomical structures, individual variability, low signal-to-noise ratio and resolution differences in medical images make automated ALD challenging.
@@
Early ALD methods rely on handcrafted image features [8, 9, 10]. While these methods achieve some success in specific applications, they depend heavily on data quality and feature design. As a result, they struggle to handle complex anatomical structures and adapt to varying imaging conditions, leading to limited robustness. With the rapid advancement of deep learning, Convolutional Neural Networks (CNNs) have transformed image processing tasks, particularly in ALD and medical image segmentation. Models such as U-Net [11] and its variants [12, 13, 14, 15, 16] are now widely applied for landmark detection. These models employ encoder-decoder architectures to effectively integrate low-level features with high-level semantic information.
@@
Transformer architectures, with their self-attention mechanisms, are particularly effective at modeling global dependencies and capturing long-range spatial information. In the field of medical imaging, Transformers [17, 18] have been increasingly integrated into image analysis tasks, often in combination with convolution-based U-Net models such as UTNet [19], TransUNet [20], and nnFormer [21]. Since accurate ALD relies heavily on global context, incorporating Transformers to capture this information has been shown to enhance detection accuracy. Notable examples include SpineHRformer [22] and DATR [23].
@@
To leverage the complementary strengths of Convolutional Neural Networks (CNNs) and Transformers, hybrid architectures such as HTC [24] and CephalFormer [25] have emerged. These models combine CNNs for local feature extraction and Transformers for modeling global dependencies, achieving significant success in ALD. However, the computational complexity of standard self-attention mechanisms scales quadratically with the number of tokens, which limits the computational efficiency and overall performance of the model to some extent. Some methods, such as Gated Axial UNet (MedT) [26], have introduced sparsity to alleviate the computational burden, but they rely on manually designed, fixed sparse patterns. These fixed patterns limit the model’s adaptability when processing different queries, as they cannot dynamically adjust the sparse pattern according to the specific context. Medical images are typically characterized by a high complexity, a low signal-to-noise ratio, and a significant domain gap between datasets from different anatomical regions and acquisition protocols. These characteristics demand models that are able to flexibly select and adjust computation regions to address these challenges. Therefore, developing a sparse attention mechanism that dynamically adjusts with queries tailored for medical image landmark detection tasks remains one of the key research directions to improve detection accuracy.
@@
To address the above challenges and forge performant ALD with accuracy and robustness, we propose a novel HYbrid ATTention Network (HYATT-Net), which incorporates a dynamic sparse attention into a hybrid Transformer-CNN architecture, specifically designed for ALD. Our approach leverages an Attention Residual Module (ARM) and a BiFormer module, marking the first combination in ALD to achieve state-of-the-art (SOTA) results. Specifically, the ARM enhances feature selection through channel and spatial attention mechanisms, while the BiFormer module leverages bi-level routing attention to effectively model long-range spatial dependencies with optimized computational and memory efficiency. Additionally, we design a Feature Fusion Correction Module (FFCM) to integrate global and local features, further improving detection accuracy and adaptability to diverse anatomical structures and imaging conditions. Deep supervision is also employed to refine predictions at multiple levels, enhancing robustness and performance.
@@
We conduct extensive experiments across five public datasets, encompassing various anatomical regions, imaging methods, and resolutions, consistently achieving state-of-the-art (SOTA) results. For example, on the ISBI2015 dataset, our method achieves a mean residual error (MRE) of 1.13mm, improving by 5% compared to HTC [24] (1.19mm) and 11% compared to FARNet [16] (1.27mm). Additionally, the Success Detection Rate (SDR) at 2mm reaches 84.78%, surpassing HTC by 1.54% and FARNet by 2.27%. Similar improvements are observed across all datasets, demonstrating the robustness, adaptability, and clear superiority of our method in high-resolution medical image analysis and landmark detection tasks.
@@
  • We propose a novel architecture that first integrates the newly designed BiFormer module with the Attention Residual Module (ARM) to enhance feature selectivity and capture global dependencies, thereby improving the accuracy and robustness of Anatomical landmark detection (ALD).
  • We incorporate deep supervision during training, along with the Feature Fusion Correction Module (FFCM), to effectively integrate global and local information, further enhancing adaptability to diverse anatomical regions and imaging modalities.
@@ 2 Related Work
2.1 Traditional ALD Approaches
@@
Although these methods achieve some success in ALD, they gradually fall short due to the high complexity and precision requirements of medical imaging. With the rapid development of deep learning, researchers begin exploring new approaches to overcome the limitations of traditional methods, significantly improving the performance and adaptability of landmark detection.
2.2 CNN or Transformer Based ALD Methods
With the rapid advancements in deep learning, Convolutional Neural Networks (CNNs) have been widely adopted for ALD. For instance, O’Neil et al. [14] propose a two-stage method based on Fully Convolutional Networks (FCN). Payer et al. [15] introduce Spatial Configuration Network (SCN), which combines local appearance features with spatial configuration information. Zhu et al. [36] propose GU2Net, a general model for multi-domain landmark detection, which improves robustness through cross-domain learning. Additionally, Ao et al. [16] develop FARNet, which incorporates multi-scale feature aggregation and refinement modules, further improving detection precision.
@@ 2.3 CNN-Transformer Hybrid Models
However, due to the complexity and low signal-to-noise ratio of medical images, high-performance models are necessary. The computational complexity of standard Transformer models becomes a bottleneck, especially when processing high-resolution medical images. Some studies, such as Swin-Unet [38] and Gated Axial UNet (MedT) [26], have incorporated sparsity into their models to address this challenge. Swin-CE [39] integrates Swin Transformer [40] encoders with convolutional encoders for cephalometric landmark detection, but these methods often rely on manually designed sparse patterns, which lack flexibility in adapting to dynamic queries.
Recent advancements in vision Transformers, particularly those utilizing dynamic sparsity, provide promising solutions. For instance, dynamic token sparsity mechanisms prune a large number of non-informative tokens, thereby accelerating the model while maintaining accuracy. In [41], dynamic token sparsity is applied to reduce computational load by removing non-informative tokens without sacrificing precision. The application of dynamic and query-aware sparse attention mechanisms for landmark detection in medical images presents a significant research opportunity. In this study, we propose employing a bi-level routing attention as a foundational dynamic sparse block, constructing a U-shaped hybrid CNN-Transformer encoder-decoder architecture for efficient landmark detection in medical images.
@@ 3 Proposed Methods
Figure 1: The overview of proposed Hybrid Attention Network(HYATT-Net). BiFormer is a module based on bilevel routing attention, and ARM stands for a Attention Residual Block. Further details will be discussed later. $N$ denote the number of landmarks.
@@ 3.1 Problem Definition
The task of ALD aims to accurately predict the locations of multiple anatomical landmarks from input medical images. Formally, given an input image $I\in\mathbb{R}^{H\times W\times C}$, the objective is to predict a set of landmarks $L=\{l_{1},l_{2},\dots,l_{N}\}$, where each $l_{i}=(x_{i},y_{i})$ represents the position of a landmark in the image. $N$ represents the number of landmarks in the image, which is also the number of output heatmap channels. The corresponding heatmap $H_{i}(x_{i},y_{i})$ for each landmark $l_{i}$ is generated using a Gaussian function:
$H_{i}=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma^{2}}\right),$ (1)
where $C$ represents the number of image channels, with $C=1$ for X-ray images, $H$ and $W$ represent the image height and width, respectively, and $\sigma$ represents the standard deviation of the Gaussian function used to generate the heatmap.
@@ 3.2 BiFormer Module Design
The BiFormer module, based on a Bi-Level Routing Attention (BRA) mechanism [41], is a key component of the proposed architecture. BRA is a dynamic, query-aware sparse attention mechanism designed to prune the least relevant key-value patchs at a coarse level, retaining only the most relevant parts for fine-grained token-level attention. Token-to-token attention is then computed within these selected patches. Compared to static sparse attention mechanisms [38, 42], BRA is more flexible in modeling long-range dependencies, making it particularly effective for dense prediction tasks that require capturing global features.
@@
Furthermore, as shown in Fig. 3, to compute the token-to-token attention for a patch $i$, initially, collect the top-$k$ most relevant patch index from matrix $I^{p}$ of patch $i$; subsequently, apply the gather algorithm to extract the corresponding $K^{g}$ and $V^{g}$ values of the relevant patches, where $K^{g},V^{g}\in\mathbb{R}^{S^{2}\times k\frac{HW}{S^{2}}\times C}$. Finally, calculate the attention between patch $i$ and the top-$k$ most relevant patches, incorporating the Local Context Enhancement (LCE), to obtain the final output $O$ for this patch. The above process can be formulated as follows:
@@
Figure 2: (a) Architecture of the BiFormer Block. (b) Architecture of the proposed Attention Residual Block. (c) Overview of the Convolutional Block Attention Module (CBAM).
@@ 3.3 Overal Network Architecture
As demonstrated in Fig. 1, the overall HYATT-Net employs a U-shaped architecture, with the main backbone consisting of hybrid Transformer-CNN modules, upsampling layers, and residual connections. The hybrid Transformer-CNN module is composed of BiFormer and ARM connected in series. The global features extracted by BiFormer are fed into ARM, which further refines local features, achieving an effective integration of global and local features. Specifically, the global information provided by BiFormer helps the model understand the geometric relationships between landmarks, while ARM refines local features based on these global features, making landmark detection more precise.
Compared to the original CNN, ARM is integrated with an attention module Convolutional Block Attention Module(CBAM) [43] to select representative features. As demonstrated in Fig. 2(c), the CBAM module combines channel attention and spatial attention, allowing the network to focus more effectively on task-relevant features, enhancing the effectiveness of feature representation and the overall performance of the network. The role of CBAM can be formulated as:
$F_{\text{out}}=\text{SA}(\text{CA}(F_{\text{in}})\odot F_{\text{in}})\odot(\text{CA}(F_{\text{in}})\odot F_{\text{in}}),$
where CA represents channel attention, SA represents spatial attention, and $\odot$ denotes element-wise multiplication.
As for input features, ARM first processes them using a dilated convolution with dilation set to 1, followed by normalization and activation function processing. Then, the processed features are passed through another dilated convolution with dilation set to 2. After normalization, the features are passed through the CBAM, and finally, the output is obtained through a residual connection. The above process illustrated in Fig. 2(b) and can be represented as:
@@
BiFormer provides rich contextual information to ARM, reducing redundant features and information loss, thereby accelerating the training process and improving convergence speed. Therefore, combining BiFormer with ARM not only enhances the model’s performance in landmark detection but also improves training efficiency and model robustness, making it suitable for complex medical images and other scenarios that require precise modeling.
As Fig. 1 shows, the downsampling part of the model consists of five layers of hybrid Transformer-CNN modules. For an input image, Patch embedding is first performed, followed by the BiFormer module for capturing global features and the ARM module for extracting local features. Multi-scale features $\{D_{1},D_{2},D_{3},D_{4},D_{5}\}$ are generated to enhance the model’s feature representation capability. These features are subsequently mapped to a feature space with 256 channels.
@@ 4.2 Baseline Methods
To verify the effectiveness of our proposed model, we compare it with several representative or SOTA methods, including U-Net [11], SCN [15], Cascade R-CNN [49], GU2Net [36], DATR [23], CeLDA [46], SR-UNet [50], FARNet [16], HTC [24]:
@@
  • SR-UNet[50]: Applies super-resolution networks and pyramid pooling to reduce errors and improve detection precision.